• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name hoyoung

    Log entry time 08:09:12 on March 06, 2012

    Entry number 364223

    keyword=Owl Shift Summary March 06, 2012

    Shift workers: Hoyoung Kang(SL), Mikhail Yurov(TO), Marco Pannunzio
    Carmignotto(3rd)
    
    Plan : BPM calibration, HARP scan, raster calibration, Helium optics
    
    01:10 Controlled access to disconnect BCM multiplier and adjust BPM
    02:20 Back to beam permit
    02:50 100 nA CW with BPMA 0/0
    03:00 Harp scan and BPM calibration at the beam position +4/+4
    03:30 Harp scan and BPM calibration at the beam position +4/-4
    03:50 Harp scan and BPM calibration at the beam position -4/-4
    04:10 Harp scan and BPM calibration at the beam position -4/+4, But we
    had problem due to beam trip and beam position fluctuation
    05:20 Slow rater calibration
    05:30 Beam off and tried to move target to carbon. But unable to move target
    06:00 Due to target motion fault, fast raster calibration started. Fast
    raster on with 100 nA CW beam. FR calibration for 15 different settings
    (1~7)x(1~7) mm and 0x0 mm
    06:45 Fast raster calibration finished, FR OFF and beam current 100 nA
    06:50 Helium optics data taking with 0 % delta
    07:10 Restricted access requested, As planned, there will be works in hall
    07:40 Hall survey finished, Restricted access